For the data set shown​ below, complete parts​ (a) through​ (d) below. x 3 4 5 7 8 y 5 7 6 13 14 ​(a)  Find the estimates of beta 0 and beta 1. beta 0almost equalsb 0equals negative 1.360 ​(Round to three decimal places as​ needed.) beta 1almost equalsb 1equals 1.919 ​(Round to three decimal places as​ needed.) ​(b)  Compute the standard​ error, the point estimate for sigma. s Subscript eequals 1.4929 ​(Round to four decimal places as​ needed.) ​(c)  Assuming the residuals are normally​ distributed, determine s Subscript b 1 Baseline . s Subscript b 1equals nothing ​(Round to three decimal places as​ needed.) ​(d)  Assuming the residuals are normally​ distributed, test Upper H 0 : beta 1 equals 0 versus Upper H 1 : beta 1 not equals 0 at the alpha equals 0.05 level of significance. Use the​ P-value approach. The​ P-value for this test is nothing. ​(Round to three decimal places as​ needed.) Make a statement regarding the null hypothesis and draw a conclusion for this test. Choose the correct answer below.
Answer #1

a)

bo=-1.361

b1=1.919

b)

SSE =Syy-(Sxy)2/Sxx= 6.686
s2 =SSE/(n-2)= 2.2287
std error σ              = =se =√s2= 1.4929

c)

estimated std error of slope =se(β1) =s/√Sxx= 0.3600

d)

test stat t = β1/se(β1)= = 5.330
p value: = 0.013

reject the null hypothesis and conclude that there is a relationship between x and y.